# The Atacama Cosmology Telescope: Quantifying Atmospheric Emission above Cerro Toco

### v1

@@ A.1 Derivation of the Matérn covariance function for scale-invariant turbulent distributions
$C(r)=\sigma^{2}\frac{2^{1-\nu}}{\Gamma(\nu)}(r\kappa_{0})^{\nu}K_{\nu}(r\kappa_{0})\quad{\rm and}\quad\Phi(\kappa)=\frac{\Gamma(\nu+\frac{3}{2})}{\pi^{\frac{3}{2}}\Gamma(\nu)}\frac{\sigma^{2}\kappa_{0}^{2\nu}}{(\kappa_{0}^{2}+\kappa^{2})^{\nu+\frac{3}{2}}}.$ (A5)
@@
$C(\tau)=\frac{2^{1-\nu}}{\Gamma(\nu)}\sigma^{2}(\tau\omega_{0})^{\nu}K_{\nu}(\tau\omega_{0})\quad{\rm and}\quad S(\omega)=\sigma^{2}\frac{\Gamma(\nu+\frac{1}{2})}{\pi^{\frac{1}{2}}\Gamma(\nu)}\frac{\tau_{0}}{(1+\omega^{2}\tau_{0}^{2})^{\nu+\frac{1}{2}}}.$ (A6)
@@ B.2 Embedding the data
Fitting over the raw ACT data simultaneously is intractable due to memory limitations. We can, however, condense the data as an “embedding” into an abstract vector space that preserves only the information necessary to fit Equation 15, i.e., the correlations between the rows of $\delta\mathbf{M}$. Decomposing $\delta\mathbf{P}$ using an SVD and keeping only the first 4 components gives us
@@ B.3 Fitting
where $\epsilon=|\mathbf{A}-\mathbf{G}\mathbf{X}||\mathbf{A}|^{-1}$ is the normalized residual of the reconstructed embedding. We adopt $\sigma=0.03$ as it is the median normalized residual after the model is fitted to all ACT data. This model is, in essence, a one-layer neural network that can be implemented in PyTorch and fitted by maximizing $\log\mathcal{L}$ using the backpropagation algorithm with the adaptive momentum (ADAM) optimizer. We fit the model over all TODs in parallel for each season. This produces a solution for $\mathbf{G}$ and $\mathbf{X}$. Despite the degeneracy mentioned in the body of the paper, we can glean from $\mathbf{G}$ the relative gain of each detector and from $\mathbf{X}$ the relative strength of each mode (i.e. the ratio of common fluctuations to non-common fluctuations).o

### v3

@@ A.1 Derivation of the Matérn covariance function for scale-invariant turbulent distributions
$C(r)=\sigma^{2}\frac{2^{1-\nu}}{\Gamma(\nu)}(r\kappa_{0})^{\nu}K_{\nu}(r\kappa_{0})\quad$ (A5)
and
$\Phi(\kappa)=\frac{\Gamma(\nu+\frac{3}{2})}{\pi^{\frac{3}{2}}\Gamma(\nu)}\frac{\sigma^{2}\kappa_{0}^{2\nu}}{(\kappa_{0}^{2}+\kappa^{2})^{\nu+\frac{3}{2}}}.$ (A6)
@@
$C(\tau)=\frac{2^{1-\nu}}{\Gamma(\nu)}\sigma^{2}(\tau\omega_{0})^{\nu}K_{\nu}(\tau\omega_{0})\quad$ (A7)
and
$S(\omega)=\sigma^{2}\frac{\Gamma(\nu+\frac{1}{2})}{\pi^{\frac{1}{2}}\Gamma(\nu)}\frac{\tau_{0}}{(1+\omega^{2}\tau_{0}^{2})^{\nu+\frac{1}{2}}}.$ (A8)
@@ B.2 Embedding the data
Fitting over the raw ACT data simultaneously is intractable due to memory limitations. We can, however, condense the data as an “embedding” into an abstract vector space that preserves only the information necessary to fit Equation 15, i.e., the correlations between the rows of $\delta\mathbf{M}$. Decomposing $\delta\mathbf{P}$ using a singular value decomposition and keeping only the first 4 components gives us
@@ B.3 Fitting
where $\epsilon=|\mathbf{A}-\mathbf{G}\mathbf{X}||\mathbf{A}|^{-1}$ is the normalized residual of the reconstructed embedding. We adopt $\sigma=0.03$ as it is the median normalized residual after the model is fitted to all ACT data. This model is, in essence, a one-layer neural network that can be implemented in PyTorch and fitted by maximizing $\log\mathcal{L}$ using the backpropagation algorithm with the adaptive momentum (ADAM) optimizer. We fit the model over all TODs in parallel for each season. This produces a solution for $\mathbf{G}$ and $\mathbf{X}$. Despite the degeneracy mentioned in the body of the paper, we can glean from $\mathbf{G}$ the relative gain of each detector and from $\mathbf{X}$ the relative strength of each mode (i.e. the ratio of common fluctuations to non-common fluctuations).